# A Benchmark for the Application of Distributed Control Techniques to the Electricity Network of the European Economic Area

### v1

@@ 1.1 The Origin of the Benchmark: Motivation and Challenges
The European Economic Area Electricity Network Benchmark (EEA-ENB) is a benchmark designed for the implementation and testing of distributed control strategies for large-scale power networks. The idea behind the benchmark is to build an abstract model of the European network of transmission systems for electricity. We represent each country of the European economic area as an independent electrical area connected to others through tie lines according to a predefined electricity network topology. The result is a real-world oriented benchmark that accounts for the presence of renewable generation and Energy Storage Systems (ESSs) in the load-frequency control (LFC) problem of the power network. The development of the EEA-ENB is essential because no established control model for the European electricity transmission system consistently serves as a reference for distributed control techniques, especially with energy storage systems and renewable energy sources. Additionally, the use case for the EEA-ENB is not restricted only to the pure development of control strategies. With minimal modifications, it can also be used for other applications, such as the economic optimization of network operation, the study of network expansion strategies, testing of security and privacy features, and simulation of emergency situations such as cascading blackouts and network restoration.
To assess the time and computation requirements for the implementation of a distributed control strategy we implement centralized Model Predictive Control (MPC) on the network. Together with the value of the cost function of centralized MPC developed, this provides the user metrics to evaluate the advantages and disadvantages in the implementation of a specific distributed control technique. The EEA-ENB is formulated with a modular approach such that extensions can be implemented if needed, allowing for various application scenarios as mentioned before. The stability of the network is assessed through the study of the Load Frequency Control (LFC) problem. Moreover, another application that is particularly relevant for this benchmark is the economical optimization of energy trading among network agents. The EEA-ENB can also be employed to formulate Distributed MPC (DMPC) techniques in the presence of hybrid dynamics thanks to a modified ESS dynamics reported. Additional extensions, not included in this work, include the characterization of each electrical area according to the deregulated energy market through the modeling of generation plants, the auction system for scheduling energy production across the various generation companies, and the market of power exchanges between different electrical areas bevrani_RobustPowerSystem_2014 .
@@ 2.1 System Description
The EEA-ENB is composed of 26 interacting electrical areas connected through tie lines and uses real-world data acquired from the ENTSO-E transparency platform accessible from _ENTSOE . Each area represents an equivalent electrical machine aggregating the inertia and dispatchable capacity of generators in that specific area, a modeling technique commonly used in the context of LFC kundur_PowerSystemStability_2022 . The electrical topology of the network is derived from the grid map also provided by ENTSO-E _ENTSOE . The benchmark is constituted by 26 control areas due to considerations about the availability and scale of the data about the 31 members of the EEA. The electrical topology of the resulting network is reported in Figure 1, where each country is labeled with the respective ISO code. Positions of the areas in the space are selected according to their geographical centroids, and, on this basis, the lengths of the tie lines are defined using the Euclidean distance as reported in Table 1.
@@
In this graph representation, each node is associated with a dynamics incorporating generation, storage, consumption, and interaction behaviors of the considered electrical area and of its neighborhood. In particular, an electrical area is composed of a multiplicity of autonomous subsystems working together to guarantee the satisfaction of the set points assigned by the area-level controller. The aggregation of those subsystems allows one to define an equivalent electrical machine for each area. Specifically, each area may comprehend the following:
@@
  • An Energy Storage System (ESS) used to accumulate and supply energy at the best convenience and according to the control strategy implemented. In general, energy storage systems can be classified into three macro-categories: electrical storage (e.g. ultracapacitors), electrochemical storage (e.g. batteries), and mechanical storage (e.g. water reservoirs). However, this distinction is not considered in the benchmark, but it is suggested as a possible extension. Following the same approach used for dispatchable generation, we consider the aggregated storage and power of all the ESSs in the electrical area.
@@ 2.2 System Dynamics
The topology of the power system is represented as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where each node $v_{i}\in\mathcal{V}$ is associated with an independent electrical area $i$, and each undirected edge $e_{ij}=e_{ji}=(v_{i},v_{j})\in\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is a tie line connecting adjacent areas $i$ and $j$, allowing for bidirectional power flow. In our case, we have 26 nodes, one for each electrical area. The presence of an edge represents the existence of a power connection. For each node $v_{i}$, we define its neighborhood as $\mathcal{N}_{i}=\{v_{j}\in\mathcal{V}\,|\,(v_{i},v_{j})\in\mathcal{E}\}$, i.e. the set of nodes connected to the node $v_{i}$. To each node $v_{i}$, $i=1,\ldots,26$ an equivalent electrical machine is associated according to the schematic in Figure 2. Each electrical area $i$ is always characterized by at least three states: the angle $\delta_{i}$ of the rotor, the operating frequency $f_{i}$ of the equivalent machine, and the energy $e_{i}$ stored in the ESS. The control inputs for the $i$-th area are the deviation $\Delta P_{i}^{\text{disp}}$ of dispatchable power production w.r.t. the scheduled value, and the power $P_{i}^{\text{ESS, c}}$ supplied to or $P_{i}^{\text{ESS, d}}$ withdrawn from the ESS. Additionally, each area is subjected to the influence of external inputs: the variation in the load request $\Delta P_{i}^{\text{load}}$, renewable energy production $\Delta P_{i}^{\text{ren}}$, and the power transmitted over the tie lines $\Delta P_{i}^{\text{tie}}$ connected to area $i$.
@@
For this system, it is common to assume bevrani_RobustPowerSystem_2014 ; kundur_PowerSystemStability_2022 a linearized discrete-time model around an operating point $(\delta_{0},f_{0})$ for the power angle and frequency dynamics. We assume $\delta_{0}=30$ [deg], but this depends on the scheduled power exchanges among electrical areas as specified later, with limits $\delta_{0}\in(0,90)$ [deg]; moreover, the operating frequency of the European power network is $f_{0}=50$ [Hz] kundur_PowerSystemStability_2022 . Regarding the ESS of the $i$-th area, the simplest model capturing the charging and discharging characteristics of a storage system is the linear representation also reported in sioshansi_EnergyStorageModelingStateoftheArt_2022 . Extensions of this model, and alternative formulations of the ESS dynamics, as the PWA description in parisio_ModelPredictiveControl_2014b , are discussed in Section 2.4. For the tie lines interaction, we also use a linearized equation kundur_PowerSystemStability_2022 under the assumption that machine angle deviations are small enough, which will be guaranteed through operating constraints in the control formulation.
@@ 2.3 Assumptions and Operating Conditions
The electrical angle deviation is constrained as $-30\leq\Delta\delta_{i}\leq 30$, so that the electrical angle satisfies $0\leq\delta_{i}\leq 60$, with $\delta_{i,0}=30$ [deg]. For the operating frequency we assume the range $-0.04\leq\Delta f_{i}\leq 0.04$, with $f_{i,0}=50$ [Hz] bevrani_RobustPowerSystem_2014 ; kundur_PowerSystemStability_2022 . For the ESSs, we consider the maximum storage capacity to be equal to the total dispatchable capacity, i.e. $0\leq e_{i}\leq P_{i,\text{max}}^{\text{disp}}$, for each area $i$. For each electrical area $i$ the following state constraints hold:
@@
$-\frac{P_{i,\text{max}}^{\text{disp}}}{\tau}\leq\Delta P_{i}^{\text{disp}}\leq\frac{P_{i,\text{max}}^{\text{disp}}}{\tau}\qquad 0\leq P_{i}^{\text{ESS, c}},\,\,P_{i}^{\text{ESS, d}}\leq\frac{P_{i,\text{max}}^{\text{disp}}}{\tau}$ (11)
@@
An aspect not usually considered in LFC systems is the total dispatchable production limit. In this benchmark, using real-world data, we want to constrain the total dispatchable production to be non-negative, and smaller than the overall capacity of a certain area, i.e. $0\leq P_{i}^{\text{disp}}(k)\leq P_{i,\text{max}}^{\text{disp}}$. Some areas may have a renewable production that exceeds the total load request. Thus, this constraint ensures that the excess is stored for later use or transmitted to neighboring areas through machine angles adjustments. As initial condition, we assume to have a dispatchable production that compensates for the total load request and that accounts for the renewable generation:
$P_{i,0}^{\text{disp}}=\max\left\{0;P_{i,0}^{\text{load}}-P_{i,0}^{\text{ren}}\right\}.$ (12)
To guarantee that $P_{i,0}^{\text{disp}}$ is not negative we select it as the maximum between zero and the difference $P_{i,0}^{\text{load}}-P_{i,0}^{\text{ren}}$.
@@ 2.4 Extensions and Alternative Formulations
Assuming that the ESSs can only be in a charging or discharging state at each time step, their dynamics can be described with the following PWA equations parisio_ModelPredictiveControl_2014b :
@@
State Augmentation and the Energy Market
The energy market can be characterized by augmenting the difference equations of system (8). Specifically, by integrating the variation of dispatchable generation and power transmission via the tie lines, and considering the electricity prices for energy generation and trade, we can effectively model the electricity market of the EEA-ENB. To this end, the dynamics (8) are augmented with the two states $P_{i}^{\text{disp}}$, $P_{i}^{\text{tie}}$, which evolve according to:
@@
In this way, we can also impose limits on the overall dispatchable generation for each electrical area according to the data acquired from _ENTSOE . Specifically, each area is subjected to $0\leq\Delta P_{i}^{\text{disp}}(k)\leq P_{i,\text{max}}^{\text{disp}}$.
@@ 2.5 Goal of the Control System
The main control goal for this system is to regulate the frequency deviation $\Delta f_{i}=f_{i}-f_{i,0}$ of each electrical area to zero, so that the frequency of the network stays at the desired value $f_{0}=50$ [Hz]. Moreover, we require to regulate the machine angle deviation $\Delta\delta_{i}=\delta_{i}-\delta_{i,0}$ to zero, such that the efficiency of the machine is preserved. In addition to the control objectives, the control system should also ensure that the operational constraints of Section 2.3 are satisfied.
However, other control goals can also be defined. First, if an economic MPC problem is formulated using electricity prices, then the total monetary cost for running the network can be considered, defined as $\text{\texteuro}/\text{MWh}$ for each agent, for each energy source, and at each time step. With this approach, the least expensive network operation strategy can be defined, trading off the lower operational cost of the network with its stability. In this regard, if soft constraints are implemented to limit the frequency deviation, then the total-time-spent and the average-time-spent outside the optimal operation interval of the frequency can also be considered as a performance indicator. Moreover, other control goals can be considered that are more specific to the technological implementation of the network. Those could regard the number of charging and discharging cycles of the batteries, their average charge level, or also the electrical machine angle deviation w.r.t. the most efficient one. This means that we might incorporate operational and maintenance costs in the benchmark and consider them as a way to compare control strategies.
@@ 3.1 Input Data
The network of equivalent machines is modeled using real data about load requests, renewable generation, and dispatchable capacities of the 26 European states selected for the implementation. Data is acquired from the ENTSO-E electricity transparency platform _ENTSOE . As an example, data for the 24 hours of January 1, 2022, are reported in Figure 3.
Refer to caption Refer to caption Refer to caption Refer to caption
@@ 3.2 Implementation Details
We use linear interpolation to compute the external inputs $\Delta P_{i}^{\text{load, meas}}(k)$ , for $k=0,\ldots,34559$, from the data of $P_{i}^{\text{load, meas}}(h)$ available every hour, for $h=1,\ldots,24$. The same approach is used for renewable generation measurements $P_{i}^{\text{ren, meas}}$, and for the forecasts $P_{i}^{\text{ren, for}}$, $P_{i}^{\text{ren, for}}$. The resulting signal variations are in Figure 4.
Refer to caption Refer to caption Refer to caption Refer to caption
@@ 3.4 Performance Metrics
To evaluate the system performance we consider a quadratic cost function (we will formally introduce and define the cost function in (17), Section 5). Specifically, since the EEA-ENB is structured for the implementation of distributed control techniques, the overall cost over one day of simulation for the network controlled with a centralized MPC architecture represents the optimality target for every alternative control formulation.
Another performance indicator of the control strategy is the computation time required to obtain the control action. Specifically, centralized MPC might not be suited for real-time control of the EEA-ENB due to the excessive computation time required to obtain the optimal control action. Distributed architectures are usually faster in obtaining the control law since they distribute the computational burden among the control agents, but they are more complex to implement. Also, the cost of communication among the control agents, represented by the total volume of information exchanged across the network, can be used as a metric.

### v3

@@ 1.1 The Origin of the Benchmark: Motivation and Challenges
The European Economic Area Electricity Network Benchmark (EEA-ENB) is a benchmark designed for the implementation and testing of distributed control strategies for large-scale power networks. The idea behind the benchmark is to build an abstract model of the European network of transmission systems for electricity. We represent each country of the European economic area as an independent electrical area connected to others through tie lines according to a predefined electricity network topology. The result is a real-world oriented benchmark that accounts for the presence of renewable generation and Energy Storage Systems (ESSs) in the Load-Frequency Control (LFC) problem of the power network.
The development of the EEA-ENB is essential because no established control model for the European electricity transmission system consistently serves as a reference for distributed control techniques, especially with energy storage systems and renewable energy sources. Additionally, the use case for the EEA-ENB is not restricted only to the pure development of control strategies. With minimal modifications, it can also be used for other applications, such as the economic optimization of network operation, the study of network expansion strategies, testing of security and privacy features, and simulation of emergency situations such as cascading blackouts and network restoration.
To assess the time and computation requirements for the implementation of a distributed control strategy we implement centralized Model Predictive Control (MPC) on the network. Together with the value of the cost function of centralized MPC developed, this provides the user metrics to evaluate the advantages and disadvantages in the implementation of a specific distributed control technique. The EEA-ENB is formulated with a modular approach such that extensions can be implemented if needed, allowing for various application scenarios as mentioned before. The stability of the network is assessed through the study of LFC problem. Moreover, another application that is particularly relevant for this benchmark is the economical optimization of energy trading among network agents. The EEA-ENB can also be employed to formulate Distributed MPC (DMPC) techniques in the presence of hybrid dynamics thanks to a modified ESS dynamics reported. Additional extensions, not included in this work, include the characterization of each electrical area according to the deregulated energy market through the modeling of generation plants, the auction system for scheduling energy production across the various generation companies, and the market of power exchanges between different electrical areas bevrani_RobustPowerSystem_2014 .
@@ 2.1 System Description
The EEA-ENB is composed of 26 interacting electrical areas connected through tie lines and uses real-world data acquired from the European Network of Transmission System Operators for Electricity (ENTSO-E) transparency platform accessible from _ENTSOE . Each area represents an equivalent electrical machine aggregating the inertia and dispatchable capacity of generators in that specific area, a modeling technique commonly used in the context of LFC kundur_PowerSystemStability_2022 . The electrical topology of the network is derived from the grid map also provided by ENTSO-E _ENTSOE . The benchmark is constituted by 26 control areas due to considerations about the availability and scale of the data about the 31 members of the EEA. The electrical topology of the resulting network is reported in Fig. 1, where each country is labeled with the respective ISO code. Positions of the areas in the space are selected according to their geographical centroids, and, on this basis, the lengths of the tie lines are defined using the Euclidean distance as reported in Table 1.
@@
In this graph representation, each node is associated with dynamics incorporating generation, storage, consumption, and interaction behaviors of the considered electrical area and of its neighborhood. In particular, an electrical area is composed of a multiplicity of autonomous subsystems working together to guarantee the satisfaction of the setpoints assigned by the area-level controller. The aggregation of those subsystems allows one to define an equivalent electrical machine for each area. Specifically, each area may comprehend the following:
@@
  • An ESS used to accumulate and supply energy at the best convenience and according to the control strategy implemented. In general, energy storage systems can be classified into three macro-categories: electrical storage (e.g. ultracapacitors), electrochemical storage (e.g. batteries), and mechanical storage (e.g. water reservoirs). However, this distinction is not considered in the benchmark, but it is suggested as a possible extension. Following the same approach used for dispatchable generation, we consider the aggregated storage and power of all the ESSs in the electrical area.
@@ 2.2 System Dynamics
The topology of the power system is represented as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where each node $v_{i}\in\mathcal{V}$ is associated with an independent electrical area $i$, and each undirected edge $\epsilon_{ij}=\epsilon_{ji}=(v_{i},v_{j})\in\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is a tie line connecting adjacent areas $i$ and $j$, allowing for bidirectional power flow. In our case, we have 26 nodes, one for each electrical area. The presence of an edge represents the existence of a power connection. For each node $v_{i}$, we define its neighborhood as $\mathcal{N}_{i}=\{v_{j}\in\mathcal{V}\,|\,(v_{i},v_{j})\in\mathcal{E}\}$, i.e. the set of nodes connected to the node $v_{i}$. To each node $v_{i}$, $i=1,\ldots,26$ an equivalent electrical machine is associated according to the schematic in Fig. 2. Each electrical area $i$ is always characterized by at least three states: the angle $\delta_{i}$ [deg] of the rotor, the operating frequency $f_{i}$ [Hz] of the equivalent machine, and the energy $e_{i}$ [GWh] stored in the ESS. The control inputs for the $i$-th area are the deviation $\Delta P_{i}^{\text{disp}}$ [GW] of dispatchable power production w.r.t. the scheduled value, and the power $P_{i}^{\text{ESS, c}}$ [GW] supplied to or $P_{i}^{\text{ESS, d}}$ [GW] withdrawn from the ESS. Additionally, each area is subjected to the influence of external inputs: the variation in the load request $\Delta P_{i}^{\text{load}}$ [GW], renewable energy production $\Delta P_{i}^{\text{ren}}$ [GW], and the power transmitted over the tie lines $\Delta P_{i}^{\text{tie}}$ [GW] connected to area $i$.
@@
For this system, it is common to assume bevrani_RobustPowerSystem_2014 ; kundur_PowerSystemStability_2022 a linearized discrete-time model around an operating point $(\delta_{0,i},f_{0,i})$ for the power angle and frequency dynamics for each area $i$. For all $i\in\{1,\ldots,26\}$, we assume $\delta_{0,i}=30$ [deg], but this depends on the scheduled power exchanges among electrical areas as specified later, with limits $\delta_{0}\in(0,90)$ [deg]; moreover, the operating frequency of the European power network is $f_{0,i}=50$ [Hz] kundur_PowerSystemStability_2022 . Regarding the ESS of the $i$-th area, the simplest model capturing the charging and discharging characteristics of a storage system is the linear representation also reported in sioshansi_EnergyStorageModelingStateoftheArt_2022 . Extensions of this model, and alternative formulations of the ESS dynamics, as the PWA description in parisio_ModelPredictiveControl_2014b , are discussed in Section 2.4. For the tie lines interaction, we also use a linearized equation kundur_PowerSystemStability_2022 under the assumption that machine angle deviations are small enough, which will be guaranteed through operating constraints in the control formulation.
@@ 2.3 Assumptions and Operating Conditions
The electrical angle deviation is constrained as $-30\leq\Delta\delta_{i}\leq 30$, so that the electrical angle satisfies $0\leq\delta_{i}\leq 60$, with $\delta_{0,i}=30$ [deg]. For the operating frequency we assume the range $-0.04\leq\Delta f_{i}\leq 0.04$, with $f_{0,i}=50$ [Hz] bevrani_RobustPowerSystem_2014 ; kundur_PowerSystemStability_2022 . For the ESSs, we consider the maximum storage capacity to be equal to the total dispatchable capacity, i.e. $0\leq e_{i}\leq e_{i,\text{max}}^{\text{disp}}$, for each area $i$, with $e_{i,\text{max}}^{\text{disp}}=1\cdot P_{i,\text{max}}^{\text{disp}}$ [GWh]. For each electrical area $i$ the following state constraints hold:
@@
$-\frac{P_{i,\text{max}}^{\text{disp}}}{1440}\leq\Delta P_{i}^{\text{disp}}\leq\frac{P_{i,\text{max}}^{\text{disp}}}{1440}\qquad 0\leq P_{i}^{\text{ESS, c}},\,\,P_{i}^{\text{ESS, d}}\leq\frac{P_{i,\text{max}}^{\text{disp}}}{1440}$ (6)
@@ 2.4 Extensions and Alternative Formulations
Assuming that the ESSs can only be in a charging or discharging state at each time step, their dynamics can be described with the following Piecewise Affine Linear (PWA) equations parisio_ModelPredictiveControl_2014b :
@@
State Augmentation and Total Production Constraints
An aspect not usually considered in LFC systems is the total dispatchable production limit. In this benchmark, using real-world data, we want to constrain the total dispatchable production to be non-negative, and smaller than the overall capacity of a certain area. To this end, the dynamics (3) are augmented with the two states $P_{i}^{\text{disp}}$, $P_{i}^{\text{tie}}$, which evolve according to:
@@
In this way, we can also impose limits on the overall dispatchable generation for each electrical area according to the data acquired from _ENTSOE . Specifically, each area is subjected to $0\leq P_{i}^{\text{disp}}(k)\leq P_{i,\text{max}}^{\text{disp}}$. Some areas may have a renewable production that exceeds the total load request. Thus, this constraint ensures that the excess is stored for later use or transmitted to neighboring areas through machine angles adjustments. As initial condition, we assume to have a dispatchable production that compensates for the total load request and that accounts for the renewable generation:
$P_{i}^{\text{disp}}(0)=\max\left\{0;P_{i}^{\text{load}}(0)-P_{i}^{\text{ren}}(0)\right\}.$ (11)
This choice of $P_{i}^{\text{disp}}(0)$ is to guarantee it to be positive, thus we select it as the maximum between zero and the difference $P_{i}^{\text{load}}(0)-P_{i}^{\text{ren}}(0)$.
@@ 2.5 Goal of the Control System
The main control goals for the benchmark are the following:
  • Regulation of the frequency deviation $\Delta f_{i}=f_{i}-f_{i,0}$ of each electrical area to zero, so that the frequency of the network stays at the desired value $f_{0}=50$ [Hz]. This is also the main goal of the control system. Moreover, we require to regulate the machine angle deviation $\Delta\delta_{i}=\delta_{i}-\delta_{i,0}$ to zero, such that the efficiency of the machine is preserved.
  • Operational constraints satisfaction. In addition to the regulation goals, the control system should also ensure that the operational constraints of Section 2.3 and 2.4 are satisfied. Ensuring that these constraints are satisfied is of primary importance for the stability of the network, and its correct functioning. Similar constraints should be enforced also for the augmented models in Section 2.4.
  • Disturbance rejection. From a control perspective, the external signals of load and renewable generation variations in (3) can be interpreted as disturbances to reject. Moreover, in the benchmark both measurements and forecasts for these signals are provided, allowing the user to represent different operational scenarios.
  • Minimization of the control effort. The control inputs of the benchmark are the variation in dispatchable generation and power exchange with the ESS for each area $i$, namely $\mathbf{u}_{i}=[\begin{matrix}\Delta P_{i}^{\text{disp}}&P_{i}^{\text{ESS, c}}&P_{i}^{\text{ESS, d}}\end{matrix}]^{\intercal}$. Minimizing the values of these quantities while ensuring the correct functioning of the system is another relevant feature for control design. The control effort is quantified as the vector product $\mathbf{u}_{i}^{\intercal}\mathbf{u}_{i}$ for each area $i$.
Other control objectives can be designed depending on the study that wants to be conducted on the benchmark. For example, if an economic MPC problem is formulated using electricity prices, then the total monetary cost for running the network can be considered, defined as $\text{\texteuro}\cdot\text{MW}$ for each agent, for each energy source, and at each time step. With this approach, the least expensive network operation strategy can be defined, trading off the lower operational cost of the network with its stability. In this regard, if soft constraints are implemented to limit the frequency deviation, then the total-time-spent and the average-time-spent outside the optimal operation interval of the frequency can also be considered as a performance indicator. Moreover, other control goals can be considered that are more specific to the technological implementation of the network. Those could regard the number of charging and discharging cycles of the batteries, their average charge level, or also the electrical machine angle deviation w.r.t. the most efficient one. This means that we might incorporate operational and maintenance costs in the benchmark and consider them as a way to compare control strategies.
@@ 3.1 Input Data
The network of equivalent machines is modeled using real data about load requests, renewable generation, and dispatchable capacities of the 26 European states selected for the implementation. Data is acquired from the ENTSO-E electricity transparency platform _ENTSOE . As an example, data for the 24 hours of January 1, 2022, are reported in Fig. 3.
@@ 3.2 Implementation Details
We use linear interpolation to compute the external inputs $\Delta P_{i}^{\text{load, meas}}(k)$ , for $k=0,\ldots,34559$, from the data of $P_{i}^{\text{load, meas}}(h)$ available every hour, for $h=1,\ldots,24$. The same approach is used for renewable generation measurements $P_{i}^{\text{ren, meas}}$, and for the forecasts $P_{i}^{\text{ren, for}}$, $P_{i}^{\text{ren, for}}$. The resulting signal variations are in Fig. 4.
@@ 3.4 Performance Metrics
The performance of the control system are measured using the value of the quadratic cost function that will be formally introduced in the control problem (12) of Section 5.1. Since the EEA-ENB is structured for the implementation of distributed control techniques, the overall cost over one day of simulation for the network controlled with the centralized MPC architecture (12) represents the optimality target for every alternative control formulation.
Another performance indicator of the control strategy is the computation time required to obtain the control action. Specifically, centralized MPC might not be suited for real-time control of the EEA-ENB due to the excessive computation time required to obtain the optimal control action. Distributed architectures are usually faster in obtaining the control law since they distribute the computational burden among the control agents, but they are more complex to implement.
As a part of the benchmark we provide the data for a control simulation of the network using the centralized MPC scheme (12) for one day of operation of the network. Both the stage cost at each step, the overall cost for one day, and the total computation time are provided as a reference for alternative control strategies. Further details are in Section 5. The end-user can consider these indexes as provided for a direct comparison, or perform their centralized MPC simulation with a different cost function to use personalized metrics.